OpenStudy (stephanos100):

PLSSS HELP!!!! MEDALS!!!! Find all solutions to the following question. x^2-5x-6 / x+2

jimthompson5910 (jim_thompson5910):

Your problem is this right? \[\Large \frac{x^2-5x-6}{x+2}\] or is it something else?

OpenStudy (stephanos100):

Yes it is like that

jimthompson5910 (jim_thompson5910):

ok so you'll factor the numerator. Think of 2 numbers that a) multiply to -6 (last term) AND b) add to -5 (middle coefficient) what two numbers meet the conditions above?

OpenStudy (stephanos100):

(x-6)(x+1)

jimthompson5910 (jim_thompson5910):

very nice

jimthompson5910 (jim_thompson5910):

ok so we now have \[\Large \frac{(x-6)(x+1)}{x+2}\]

jimthompson5910 (jim_thompson5910):

nothing will cancel since we don't have any common factors

jimthompson5910 (jim_thompson5910):

is the original expression in your problem set equal to anything? like equal to 0?

OpenStudy (stephanos100):

yeah equal to zero

jimthompson5910 (jim_thompson5910):

so the problem is really this? \[\Large \frac{x^2-5x-6}{x+2}=0\]

OpenStudy (stephanos100):

yes

jimthompson5910 (jim_thompson5910):

ok awesome

jimthompson5910 (jim_thompson5910):

the denominator will NEVER be equal to zero. You cannot divide by zero

jimthompson5910 (jim_thompson5910):

So if the fraction equals zero, then only the numerator can be zero

jimthompson5910 (jim_thompson5910):

that means if \[\Large \frac{x^2-5x-6}{x+2}=0\] then \[\Large x^2-5x-6=0\] agreed?

OpenStudy (stephanos100):

agreed.

jimthompson5910 (jim_thompson5910):

recall that you factored x^2-5x-6 to get (x-6)(x+1)

jimthompson5910 (jim_thompson5910):

So \[\Large x^2-5x-6=0\] turns into \[\Large (x-6)(x+1)=0\]

OpenStudy (stephanos100):

yes so then the zeros would be -1 and 6

jimthompson5910 (jim_thompson5910):

from here you use the zero product property

jimthompson5910 (jim_thompson5910):

`yes so then the zeros would be -1 and 6` very nice

jimthompson5910 (jim_thompson5910):

those are your two results

OpenStudy (stephanos100):

Ok thank you!!!!

jimthompson5910 (jim_thompson5910):

no problem